John is 3 years older than Sam. The product of their ages is 54. Find their ages
Answer by mananth(16949)

Sam = x
John = x+3
...
x(x+3)=54
x^2+3x=54
x^2+3x-54=0
x^2+9x-6x-54=0
x(x+9)-6(x+9)=0
(x-6)(x+9)=0
x= 6 OR x =-9
(6 & 9) OR (-9, -6)
,,
m.ananth@hotmail.ca